OpenStudy (anonymous):

Simplify the given expression. √-49 over (3+4i)-(2-5i) Answers: 63+7i over 82 -63+7i over 82 7+7i over 2 -7+7i over 2

OpenStudy (lalaly):

7i / 1+9i 7i-63/ 82

OpenStudy (lalaly):

the second 1

OpenStudy (angela210793):

7i/(1+9i)=7i(1-9i)/(1+9i)*(i-9i)=(7i-63)/82=-63+7i over 82